# Born a Transformer – Always a Transformer?

### v2

@@ 1 Introduction
Transformers (Vaswani et al., 2017) have become the backbone of most LLMs (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023). Given their importance, there has been a long line of research to understand their theoretical underpinnings. Early analyses showed that transformers can approximate any sequence-to-sequence function at fixed input length (Yun et al., 2019) and simulate Turing machines (Pérez et al., 2019) when given unbounded compute. However, subsequent studies have uncovered fundamental expressibility and learnability limitations, finding classes of problems which transformers cannot express across arbitrary input lengths (e.g. Hahn, 2020; Strobl et al., 2024; Merrill and Sabharwal, 2023; Sanford et al., 2024), or which transformers systematically struggle to learn and generalize on (e.g. Hahn and Rofin, 2024; Zhou et al., 2024). Empirical work confirms that small transformers trained from scratch do instantiate such limitations (e.g. Bhattamishra et al., 2020; Huang et al., 2025; Delétang et al., 2023; Zhou et al., 2024). For instance, transformers trained from scratch struggle to generalize to longer or more challenging examples even on tasks as simple as copying a long string with repeated symbols (Zhou et al., 2024), or retrieving which token followed the last occurrence of some query token in a long context (Liu et al., 2024a)—findings in close agreement with theoretical arguments about transformer’s learning biases (Huang et al., 2025).
@@
However, it remains unclear if such results on expressiveness and generalization have any bearing on large-scale pretrained models. Modern LLMs have many more layers, and many orders of magnitude more parameters than the small-scale transformers typically used for testing theoretical limitations. Model scale alone might make theoretical limitations irrelevant at realistic input lengths. Perhaps even more importantly, LLMs are pretrained on a massive scale, with billions of tokens of varied data that may impart fundamentally new inductive biases and algorithmic abilities (Rytting and Wingate, 2021) that may be invisible in models trained from scratch on the target task alone (Amos et al., 2024). Indeed, LLMs’ capabilities scale in ways that are difficult to foresee: empirical “laws” relate performance to model, data, and compute budget (Kaplan et al., 2020; Hoffmann et al., 2022) and qualitatively new behaviors—few-shot in-context learning, chain-of-thought reasoning, tool use—emerge abruptly at scale (Wei et al., 2022; Ganguli et al., 2022; Bubeck et al., 2023). Since pretraining reshapes both the model’s inductive biases and the sub-circuits it relies on, length-generalization failures observed in small models might not transfer automatically to their large-scale counterparts.
@@
Scope. We focus on retrieval and copying as fundamental operations that provide a test bed for answering these questions. These are fundamental operations in the usage of LLMs—e.g., retrieving from context via induction heads is key to performing in-context learning (Olsson et al., 2022; Song et al., 2025; Crosbie and Shutova, 2025), while faithfully copying exact spans makes retrieval-augmented generation more effective (Yu et al., 2024). Understanding the potential impact of architectural limitations on LLMs is thus of great interest for these tasks. In this work, we revisit the theory of length generalization for these operations, ask how its predictions hold up for pretrained transformers, and diagnose which abilities are amplified by pretraining and which limitations persist.
@@
We study a family of retrieval and copying tasks222i.e., copying of the input, not the training data, similar to Jelassi et al. (2024) (Figure 1) inspired by Liu et al. (2024a); Zhou et al. (2024). We organize retrieval into two umbrella categories of unique and non-unique induction, respectively, in which the query token appears exactly once or multiple times respectively. In each variant, the model must return a token located either before or after a designated occurrence of that query token. We make an analogous distinction in copying between copying unique tokens or repeated tokens, and between copying in the forward or the reverse direction.
@@
All retrieval and copying tasks we consider are in in principle expressible by transformers; however, theory predicts systematic differences in length generalizability – that is, in transformers’ ability to correctly perform on instances longer than those seen during training.
Specifically, we show (Section 3) that theory predicts that unique-token induction is easier than its non-unique counterpart (Huang et al., 2025; Zhou et al., 2024); that forward and reverse variants are equally hard; and that, within non-unique induction, retrieving is easier from the first occurrence of the query character than the last one. However, given the impressive capabilities of transformer-based LLMs, it is reasonable to posit that the massive scale of both such models and of pretraining might overcome these limitations. For instance, high prevalence of non-unique copying in natural language might lead transformers to learn specialized circuits or components for handling such situations.
@@
Here, we directly test whether these architectural limitations remain relevant to LLMs. Our findings show that large-scale pretraining selectively amplifies certain abilities more than others (Section 4.1). In addition to showing our results on synthetic settings, we show how these limits might surface and the reliability risks they pose for practitioners in more natural settings as well (Section 4.2). While targeted fine-tuning can even out this selective amplification of abilities born out of pretraining, we find that fundamental architectural limits still bind pretrained transformers (Section 4.3).
@@ 2 Background, Notation & Definitions
We examine two primary task types, Retrieval and Copying, each instantiated under Unique and Non-unique conditions, as illustrated in Figure 1. In Retrieval, the sequence has the form $X=\langle bos\rangle x_{1},\dots x_{N-1}\langle sep\rangle x_{N}$, where each $x_{i}\in$ some alphabet $\Sigma$ and $x_{N}=q$ is the query token. The query $q$ appears $t$ times in the context, indexed as $q_{1},\dots q_{t}$ with $1\leq q_{1}<\dots<q_{t}\leq N$. Transformer has to predict a single token as the output given the input. For the Unique condition, we define UL (Unique Induction Left) and UR (Unique Induction Right) respectively when the output expected is the token preceding or following the query’s single occurrence, i.e., $x_{q_{1}-1}$ and $x_{q_{1}+1}$. Under the Non-unique condition, NLFirst (Non-unique Left First) and NRFirst (Non-unique Right First) use the first occurrence to predict $x_{q_{1}-1}$ and $x_{q_{1}+1}$ as the output, while NLLast (Non-unique Left Last) and NRLast [^3] The NRLast variant is conceptually equivalent to the Flip-Flop task of Liu et al. (2024a). (Non-unique Right Last) use the last occurrence to produce $x_{q_{t}-1}$ and $x_{q_{t}+1}$ respectively. For Copying, the sequence is $X=\langle bos\rangle x_{1}\dots x_{N}\langle sep\rangle$ and the expected output is a continuation $x_{N+1}\dots x_{2N}$. The decoder is expected to replicate the latter segment based on the first. For Unique conditions all of the tokens in $X$ are unique. UF (Unique Forward copying) and UB (Unique Backwards copying) denote forward and reverse copying of the input segment. For Non-unique settings there is no uniqueness constraint on the $x_{i}$ values and NF and NB also denote the forward and reverse copying of the input segment.
@@
Induction heads were characterized by Elhage et al. (2021) as part of a two-head circuit present in autoregressive transformers and were later shown to be critical to performing in-context learning (Olsson et al., 2022). A previous-token head in layer $\ell$ writes the embedding of the immediately preceding token into the residual stream. An induction head in some layer $\ell^{\prime}\!>\!\ell$ then queries for positions whose stored “previous token” equals the current token and copies the corresponding value vector forward, enabling the model to predict the token after a repeated bigram. Our UR retrieval as well as the UF copying task setups are not novel and are meant to test this induction circuit ability and compare with our other task setups.
@@ 3 Theoretical Length Generalization Guarantees for Our Tasks
We derive theoretical predictions for transformers’ length-generalization abilities on the retrieval and copying tasks introduced above, using the framework of Huang et al. (2025), which links length-generalizability to expressiveness in C-Rasp (Yang and Chiang, 2024). The Rasp programming language (Weiss et al., 2021) offers a human-readable, symbolic syntax that mirrors transformer computations, making it easier to reason about their behavior. However, Rasp is overly expressive: it permits direct positional interaction between token indices that real transformers cannot realise – motivating variants that restrict such operations (Zhou et al., 2024). Of particular interest is C-Rasp (Yang and Chiang, 2024), which restricts the usage of positional information and arithmetic operations. Huang et al. (2025) further added positionally-aware primitives to C-Rasp to construct the language C-Rasp[pos] [^4] Referred to as C-Rasp[local, periodic] in Huang et al. (2025). See Appendix A.1.. Doing so, they established–both theoretically and empirically–a tight link between the existence of a C-Rasp[pos] program for a task and the ability of decoder-only transformers (with absolute or no positional encodings) to length-generalize on that same task: in particular, if C-Rasp[pos] program exists, a decoder-only model provably generalizes to longer inputs under a specific formal model of training (Theorem 7 in Huang et al. (2025)). Conversely, extensive experimental evidence in Huang et al. (2025) suggests that, if no C-Rasp program exists, such models fail to generalize to longer inputs. [^5] While Huang et al. (2025) focuses on absolute (APE) or no positional encodings (NoPE), our empirical results show that the conclusions also match behavior of Rotary Positional encodings (RoPE) (See Appendix B). We therefore adopt the following strategy for providing length generalizing guarantees across our retrieval and copying tasks:
Construct a C-Rasp[pos] program the task length-generalizes to arbitrary $\text{\emph{Construct a {C}-{Rasp}{[pos]} program} }\Longrightarrow\text{the task length-generalizes to arbitrary $n$.}$
@@
That is, C-Rasp[pos] expressivity is the same for UL and UR, the same for UF and UB, etc. (Corollary 9). All tasks are annotated in Figure 1 with their C-Rasp[pos] expressibility.
Figure 3: In-context accuracy for Qwen2.5-32B and Llama-3 70B across all our tasks averaged over 3 seeds. The error bars show prompt sensitivity and how accuracy is affected by different prompts. Across all settings, lengths, model size, and task type, retrieving the token to the left of the query token is always more difficult compared to the one to the right, provided all other things are constant. Similarly, copying in the forward direction is easier than copying backwards. Detailed prompts, similar performance graphs on other models (including instruction tuned variants) are in Appendix C.1.
@@ 4.1 Eliciting Abilities via In-Context Learning
We evaluate pretrained completion as well as instruction–tuned variants of the following models: Llama3.1--8B, Llama3.1--70B, Qwen2.5--7B, and Qwen2.5--32B. For retrieval as well as copying, each model is evaluated across various prompt variations. The prompt variations relate to the few-shot example size (examples are equal in length to the test string or are smaller but clearer) and the instruction templates (details on prompt templates in the Appendix C.1) with the templates adjusted for completion as well as instruction-tuned models. All experiments use exactly $1,500$ test strings for each length ${10,20,30,40,50}$ generated under three independent seeds. We cap the input string length at $50$ tokens to maintain a bounded vocabulary comparable to the unique setups; consequently, the behavior of the models on substantially longer spans is still unknown. The test strings contain spaced ASCII letters and digits, to ensure that every character gets tokenized as an independent token and thus, the required uniqueness / repetitions can be maintained within the context, whenever required. Few-shot prompts contain $k=5$ demonstrations, reserving one slot whenever an explicit explanation is required by the template. For each task that belongs to the same group i.e. for Unique (UR, UL) and for Non-Unique (NRLast, NRFirst, NLLast, NLFirst), we have the same strings on which we test the in-context ability with different expected answer to avoid the confound of different datapoints.
@@
Pretraining induces an in–context asymmetry. We found a persistent left–right asymmetry in all of the pretrained models we tested (Figure 3). For each matched pair group—whether in retrieval or copying—the rightward variant consistently performs better, across every model, prompt template, and sequence length. Because this asymmetry in performance also appears when we strip all natural-language instructions and provide only input–output pairs for the model to infer the task in-context, the drop in performance cannot be attributed to lexical cues such as the words “right” or “forward” (explored in earlier work (He et al., 2025)). C-Rasp[pos] predicts symmetric performance in transformers trained from scratch within a task family (Theorem 2) and our constructions for solving these tasks involve the usage of induction and anti-induction circuits. We therefore hypothesize that pretraining instills this bias because induction is a more natural task than anti-induction and thus more induction head circuits get formed during pretraining as opposed to anti-induction circuits; Sections 4.3 and 4.4 probe this explanation in more detail.
@@
Unique tasks are easier to elicit than non-unique tasks. Unique retrieval and copying setups both perform relatively well in rightward / forward direction. In contrast, in the non unique settings, eliciting correct retrieval in any direction is challenging across task setups. Specifically, the underperformance of NRFirst in pretrained models does not align with what is expected for transformers trained from scratch (Theorem 1). Copying, by contrast, is easier to elicit than retrieval, even though copying requires predicting many tokens and hence offers more opportunities for error. For instance, Non-unique-forward (NF) remains near perfect throughout. As we only consider input strings till length 50 to have comparable results vis-a-vis unique settings, the behavior of models on longer spans is unknown. The next section therefore stress-tests in-context copying on longer paragraphs, seeking to reveal if the failure modes predicted by C-Rasp[pos] for NF appear in realistic settings.
@@
Takeaway: Pretraining introduces an asymmetry favoring retrieving and copying in the forward/right direction over the backwards/left. Pretraining also selectively amplifies abilities of a transformer, as evident from the near perfect performance on copying repeated tokens, and poor performance on certain retrieval tasks, even though the latter are better learnable in a length-generalizable way for the architecture.
4.2 In-Context Copying in Natural Settings as a Test Bed

### v3

@@ 1 Introduction
Transformers (Vaswani et al., 2017) have become the backbone of most LLMs (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023). Given the widespread usage of LLMs, there has been a long line of research to understand their theoretical underpinnings. Early analyses showed that transformers can approximate any sequence-to-sequence function at fixed input length (Yun et al., 2019) and even simulate Turing machines (Pérez et al., 2019) when given unbounded compute. However, subsequent studies have uncovered fundamental expressibility and learnability limitations, finding classes of problems which transformers cannot express across arbitrary input lengths (e.g. Hahn, 2020; Strobl et al., 2024; Merrill and Sabharwal, 2023; Sanford et al., 2024), or which transformers systematically struggle to learn and generalize on (e.g. Hahn and Rofin, 2024; Zhou et al., 2024). Empirical work confirms that small transformers trained from scratch do instantiate such limitations (e.g. Bhattamishra et al., 2020; Huang et al., 2025; Delétang et al., 2023; Zhou et al., 2024). For instance, transformers trained from scratch struggle to generalize to longer or more challenging examples even on tasks as simple as copying a long string with repeated symbols (Zhou et al., 2024), or retrieving which token followed the last occurrence of some query token in a long context (Liu et al., 2024a) — findings in close agreement with theoretical arguments about transformer’s learning biases (Huang et al., 2025).
@@
However, it remains unclear if such results on expressiveness and generalization have any bearing on large-scale pretrained models. Modern LLMs have many more layers, and many orders of magnitude more parameters than the small-scale transformers typically used for testing theoretical limitations. Model scale alone might make theoretical limitations irrelevant at realistic input lengths. Perhaps even more importantly, LLMs are pretrained on a massive scale, with trillions of tokens of varied data that may impart fundamentally new inductive biases and algorithmic abilities (Rytting and Wingate, 2021) that may be invisible in models trained from scratch on the target task alone (Amos et al., 2024). Indeed, LLMs’ capabilities scale in ways that are difficult to foresee: empirical “laws” relate performance to model, data, and compute budget (Kaplan et al., 2020; Hoffmann et al., 2022) and qualitatively new behaviors such as in-context learning, chain-of-thought reasoning, and tool use emerge abruptly at scale (Wei et al., 2022; Ganguli et al., 2022; Bubeck et al., 2023). Since pretraining reshapes both the model’s inductive biases and the sub-circuits it relies on, length-generalization failures observed in small models might not transfer automatically to their large-scale counterparts.
@@
Scope. We focus on retrieval and copying as fundamental operations that provide a test bed for answering these questions. These operations are critical building blocks for many real-world LLM applications, such as question answering and summarization (Wiegreffe et al., 2025; Fan et al., 2024). On a mechanistic level, retrieving from context via induction heads is key to performing in-context learning (Olsson et al., 2022; Song et al., 2025; Crosbie and Shutova, 2025), while faithfully copying exact spans makes retrieval-augmented generation more effective (Yu et al., 2024). Understanding the potential impact of architectural limitations on LLMs is thus of great interest for these tasks.
@@
In this work, we revisit the theory of length generalization for these operations, ask how its predictions hold up for pretrained transformers, and diagnose which abilities are amplified by pretraining and which limitations persist. We study a family of retrieval and copying tasks222i.e., copying of the input, not the training data, similar to Jelassi et al. (2024) (Figure 1) inspired by Liu et al. (2024a); Zhou et al. (2024). We organize retrieval into two umbrella categories of unique and non-unique induction, respectively, in which the query token appears exactly once or multiple times respectively. In each variant, the model must return a token located either before or after a designated occurrence of that query token. We make an analogous distinction in copying between copying unique tokens or repeated tokens, and between copying in the forward or the reverse direction.
@@
All retrieval and copying tasks we consider are expressible by transformers in principle; however, theory predicts systematic differences in length generalizability333ability to correctly solve tasks on instances longer than those seen during training. Specifically, we show (Section 3) that theory predicts transformers to have a uniqueness bias (i.e., unique-token induction is easier than its non-unique counterpart) (Huang et al., 2025; Zhou et al., 2024); but for it to not have a directional bias (i.e., that forward and reverse variants are equally hard). Theory also predicts that within non-unique induction, retrieving is easier from the first occurrence of the query character than from the last one.
@@
However, given the impressive capabilities of transformer-based LLMs, it is reasonable to posit that the massive scale of both such models and of pretraining might overcome these limitations. For instance, high prevalence of non-unique copying in natural language might lead transformers to learn specialized circuits or components for handling such situations. Here, we directly test whether these architectural limitations remain relevant to pretrained LLMs. We study 8 pretrained LLMs from Llama-3 and Qwen2.5 families (Grattafiori et al., 2024; Yang et al., 2024). Our findings show that large-scale pretraining produces a directional bias (Section 4.1). In addition to showing our results on synthetic settings, we show how these limits might surface and the reliability risks they pose for practitioners in more natural settings as well (Section 4.2). While targeted fine-tuning can even out the directional bias born out of pretraining, we find that the fundamental architectural uniqueness bias still binds pretrained transformers (Section 4.3).
@@ 2 Background, Notation & Definitions
We examine two primary task types, Retrieval and Copying, each instantiated under Unique and Non-unique conditions, as illustrated in Figure 1. In Retrieval, the sequence has the form $X=\langle bos\rangle x_{1},\dots x_{N-1}\langle sep\rangle x_{N}$, where each $x_{i}\in$ some alphabet $\Sigma$ and $x_{N}=q$ is the query token. The query $q$ appears $t$ times in the context, indexed as $q_{1},\dots q_{t}$ with $1\leq q_{1}<\dots<q_{t}\leq N$. Transformer has to predict a single token as the output given the input. For the Unique condition, we define UL (Unique Induction Left) and UR (Unique Induction Right) respectively when the output expected is the token preceding or following the query’s single occurrence, i.e., $x_{q_{1}-1}$ and $x_{q_{1}+1}$. Under the Non-unique condition, NLFirst (Non-unique Left First) and NRFirst (Non-unique Right First) use the first occurrence to predict $x_{q_{1}-1}$ and $x_{q_{1}+1}$ as the output, while NLLast (Non-unique Left Last) and NRLast [^4] The NRLast variant is conceptually equivalent to the Flip-Flop task of Liu et al. (2024a). (Non-unique Right Last) use the last occurrence to produce $x_{q_{t}-1}$ and $x_{q_{t}+1}$ respectively.
For Copying, the sequence is $X=\langle bos\rangle x_{1}\dots x_{N}\langle sep\rangle$ and the expected output is a continuation $x_{N+1}\dots x_{2N}$. The decoder is expected to replicate the latter segment based on the first. For Unique conditions all of the tokens in $X$ are unique. UF (Unique Forward copying) and UB (Unique Backwards copying) denote forward and reverse copying of the input segment. For Non-unique settings there is no uniqueness constraint on the $x_{i}$ values and NF and NB also denote the forward and reverse copying of the input segment.
@@
Induction heads were characterized by Elhage et al. (2021) as part of a two-head circuit present in autoregressive transformers
@@
and were later shown to be critical to performing in-context learning (Olsson et al., 2022).A previous-token head in layer $\ell$ writes the embedding of the immediately preceding token into the residual stream. An induction head in some layer $\ell^{\prime}\!>\!\ell$ then queries for positions whose stored “previous token” equals the current token and copies the corresponding value vector forward, enabling the model to predict the token after a repeated bigram. Our UR retrieval as well as the UF copying task setups are not novel and are meant to test this induction circuit ability and compare with our other task setups.
@@ 3 Theoretical Length Generalization Guarantees for Our Tasks
We derive theoretical predictions for transformers’ length-generalization abilities on the retrieval and copying tasks introduced above, using the framework of Huang et al. (2025), which links length-generalizability to expressiveness in C-Rasp (Yang and Chiang, 2024). Of particular interest is C-Rasp (Yang and Chiang, 2024), which restricts the usage of positional information and arithmetic operations. Huang et al. (2025) further added positionally-aware primitives to C-Rasp to construct the language C-Rasp[pos] [^5] Referred to as C-Rasp[local, periodic] in Huang et al. (2025). See Appendix A.1.. Doing so, they established–both theoretically and empirically–a tight link between the existence of a C-Rasp[pos] program for a task and the ability of decoder-only transformers (with absolute or no positional encodings) to length-generalize on that same task: in particular, if C-Rasp[pos] program exists, a decoder-only model provably generalizes to longer inputs under a specific formal model of training (Theorem 7 in Huang et al. (2025)). More precisely, this guarantee applies to APE Transformers under an idealized training procedure, where all training data up to some maximum input length $N$ are available and the exact minimizer of a regularized loss is found. Under these specific conditions, the existence of a C-Rasp[pos] program for that task guarantees length generalization to larger lengths. While theory has only been worked out for this specific setup, empirically we observe converging results across different PE types and realistic SGD-based training (as shown in Huang et al. (2025) as well as in the experiments in Appendix B). As a converse to this length generalization guarantee, extensive experimental evidence in Huang et al. (2025) suggests that, if no C-Rasp program exists, such models fail to generalize to longer inputs. [^6] While Huang et al. (2025) focuses on absolute (APE) or no positional encodings (NoPE), our empirical results show that the conclusions also match the behavior of Rotary Positional encodings (RoPE) (See Appendix B). Full theoretical understanding for RoPE length generalization remains an interesting problem. One interesting aspect is that, while training RoPE for UF and UB did not show perfect generalization, we did obtain perfect generalization when instead autoregressively decoding from models trained for UR/UL. In APE, we obtained perfect generalization when training directly for UF and UB, in line with C-Rasp[pos] expressivity.
We therefore adopt the following strategy for providing length generalizing guarantees across our retrieval and copying tasks:
[…]
@@
That is, C-Rasp[pos] expressivity is the same for UL and UR, the same for UF and UB, etc. (Corollary 9). All tasks are annotated in Figure 1 with their C-Rasp[pos] expressibility. Experiments in Appendix B show that transformers trained from scratch on these tasks exhibit length generalization as predicted by the theoretical results.
Figure 3: In-context accuracy for Llama‑3 70B and Qwen2.5‑32B across all our tasks averaged over three seeds. Across all settings, lengths, model size, and task type, we observe a Directional Bias: Retrieving the token to the left of the query token is always more difficult compared to the one to the right, provided all other things are constant. Similarly, copying in the forward direction is easier than copying backwards. Detailed prompts, similar performance graphs on other models (including instruction-tuned variants) are in Appendix C.
@@ 4.1 Eliciting Abilities via In-Context Learning
We evaluate pretrained completion as well as instruction-tuned variants of the following models: Llama3.1--8B, Llama3.1--70B, Qwen2.5--7B, and Qwen2.5--32B. For retrieval as well as copying, each model is evaluated across various prompt variations. The prompt variations relate to the few-shot example size (examples are equal in length to the test string or are smaller but clearer) and the instruction templates (details on prompt templates in the Appendix C.1) with the templates adjusted for completion as well as instruction-tuned models. All experiments use 1,500 test strings for each length 10,20,30,40,50,100{10,20,30,40,50,100} generated under three independent seeds. We cap the input string length at $100$ tokens to maintain a bounded vocabulary comparable to the unique setups; see Appendix C.4 for results at other lengths. The test strings contain spaced letters (including diacritics) and digits, to ensure that every character gets tokenized as an independent token and thus, the required uniqueness / repetitions can be maintained within the context, whenever required. Few‑shot prompts contain $k=5$ demonstrations, reserving one slot whenever an explicit explanation is required by the template. For each task that belongs to the same group i.e. for Unique (UR, UL) and for Non-Unique (NRLast, NRFirst, NLLast, NLFirst), we have the same strings on which we test the in-context ability with different expected answers to avoid the confound of different datapoints.
@@
Directional Bias: Pretraining induces a left-right asymmetry. We found a persistent directional bias in all of the pretrained models we tested (Figure 3). For each task, retrieval or copying, the models consistently performs better on rightward tasks (UR, NRFirst, NRLast, UF, NF), across models, prompt templates, and sequence lengths. The drop in performance for the leftward tasks cannot be attributed to lexical cues such as the words “right” or “forward” (explored in earlier work (He et al., 2025)) because the performance asymmetry persists even when we strip all natural‑language instructions and provide only input–output pairs for the model to infer the task in-context. C-Rasp[pos] predicts symmetric performance in transformers trained from scratch within a task family (Theorem 2). We further find that models trained from scratch do not exhibit such a left-right asymmetry, confirming it is not inherent to the architecture (Appendix B). We therefore hypothesize that pretraining instills this bias because induction is a more natural task than anti‑induction and thus more induction head circuits get formed during pretraining as opposed to anti-induction circuits; Sections 4.3 and 4.4 probe this explanation in more detail.
@@
Uniqueness Bias: Unique tasks are easier to elicit than non‑unique tasks. Unique retrieval and copying setups both perform relatively well in rightward / forward direction. In contrast, in the non-unique settings, eliciting correct retrieval in any direction is challenging across task setups. We observe an interesting difference to what is expected for transformers trained from scratch: First, the underperformance of NRFirst in pretrained models does not align with what is expected for transformers trained from scratch (Theorem 1). Copying, by contrast, is easier to elicit than retrieval, even though copying requires predicting many tokens and hence offers more opportunities for error. Remarkably, Non‑unique‑forward (NF) remains near perfect throughout. As this experiment only considers input lengths up to 100 for comparable results vis-à-vis the unique settings, an interesting question is how copying performs at longer lengths. The next section therefore stress‑tests in‑context copying on longer paragraphs, seeking to reveal if the theoretically predicted difficulty of non-unique copying (NF) appears in realistic settings.
@@
Takeaway: Pretraining introduces a Directional Bias favoring retrieving and copying in the forward/right direction over the backwards/left. Pretraining also selectively amplifies the abilities of a transformer, as evident from the near perfect performance on copying repeated tokens, and poor performance on certain retrieval tasks, even though the latter are better learnable in a length-generalizable way for the architecture.
4.2 Testing Uniqueness Bias and Directional Bias in Natural Settings
Testing Uniqueness Bias by Copying Lorem Ipsum Paragraphs.